A 0.157kg glider is moving to the right on a frictionless, horizontal air track with a speed of 0.850m/s . It has a head-on collision with a 0.306kg glider that is moving to the left with a speed of 2.26m/s . Suppose the collision is elastic.

Find the magnitude of the final velocity of the 0.157kg glider.


Find the magnitude of the final velocity of the 0.306kg glider.
Answer #1

uploaded imageuploaded image

v1 =[0.85(0.157-0.306) + (-2.26)(2*0.306)]/(0.306+0.157) = -3.26 m/s towards left

v2 = -2.26(0.306-0.157) + 2*0.157*0.85 /(0.306+0.157) = -0.15 m/s towards left